the car initially at rest undergoes uniform acceleration for 6.32 seconds and covers a distance of 120 m what is approximate acceleration of the car
1
Expert's answer
2020-10-23T11:51:15-0400

We can find the acceleration of the car from the kinematic equation:


"s=v_0t+\dfrac{1}{2}at^2,"

here, "s=120\ m" is the distance traveled by the car, "v_0=0\ \dfrac{m}{s}" is the inital velocity of the car, "a" is the acceleration of the car and "t=6.32\ s" is the time during which the car accelerates.

Then, from this equation we can find the acceleration of the car:


"a=\dfrac{2s}{t^2}=\dfrac{2\cdot 120\ m}{(6.32\ s)^2}=6.0\ \dfrac{m}{s^2}."

Answer:

"a=6.0\ \dfrac{m}{s^2}."